# Stable Differentiable Modal Synthesis for Learning Nonlinear Dynamics

### v1

@@ 7  CONCLUSION
Future work will be focused on using this framework for learning from acoustic recordings of string instruments. This would require improvements in the training methodology, most notably in the training loss, as we would not have the reference data for displacements and velocities of each mode. It is theorised that the proposed approach could be used to expand the timbral range of a real instrument in a digital domain and synthesise sounds which were not part of the recorded data by changing physical parameters after the training. In addition, physics-informed architectures like that presented here have the potential to address problems where the underlying physical phenomena are not fully understood, e.g., the bowed string in musical acoustics [10].

### v2

@@ 7  CONCLUSION
Future work will be focused on using this framework for learning from acoustic recordings of string instruments. This will introduce additional challenges. First, modal frequencies and the loss profile must be estimated from audio spectrograms because physical parameters of a string will be unknown. Second, external excitation should be formulated as an initial condition (e.g., a triangular function) since estimation of an excitation force would require additional sensors [24]. Finally, the training methodology must be robust to noise and based on audio waveforms as we will not have the reference data for displacements and velocities of each mode. It is theorised that the proposed approach could be used to expand the timbral range of a real instrument in a digital domain and synthesise sounds which were not part of the recorded data by changing physical parameters after the training. In addition, physics-informed architectures like that presented here have the potential to address problems where the underlying physical phenomena are not fully understood, e.g., the bowed string in musical acoustics [10].